# The 2-systole on compact Kähler surfaces with positive scalar curvature

### v1

@@ 3. The systolic inequality on $\mathbb{P}^{2}$ and its blow-up
$\mathcal{J}_{\mathbb{P}^{2}}([\omega])=4\pi\,\operatorname{sys}^{\operatorname{hol}}_{2}(\omega)\,\frac{c_{1}(\mathbb{P}^{2})\cdot[\omega]}{[\omega]^{2}}=4\pi\cdot a\cdot\frac{3a}{a^{2}}=12\pi.$ (3.1)
@@
for every Kähler class $\omega]$ on $X_{k}$ and every $k\geq 1$. Since $\mathcal{J}_{X_{k}}$ is defined purely in terms of intersection numbers, this is intrinsically a finite–dimensional optimization problem on the Kähler cone. However, the complexity of the Mori cone $\overline{\mathrm{NE}}(X_{k})$ increases rapidly with $k$: for large $k$ there are many extremal rays, and for $k\geq 9$ (in the non–del Pezzo regime) the cone is not even finitely generated. A direct ray–by–ray analysis of all effective curve classes is therefore hopeless.
@@
For a given Kähler class we set $m:=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$. In particular, due to each exceptional curve $E_{i}$ is effective and $\omega]\cdot E_{i}=t_{i}$, we always have $m\leq t_{i}$ for all $i$. Thus, if we fix $(a,m,S)$ and vary the individual $t_{i}$ subject to
@@
then $a$ and $S$ are held fixed while $\mathcal{J}_{X_{k}}([\omega])$ changes only through the quantity $Q$ appearing in $\omega]^{2}$. In particular, for fixed $(a,m,S)$ the maximal possible value of $\mathcal{J}_{X_{k}}([\omega])$ is attained when $Q$ is as large as allowed by these constraints. This leads to the following elementary “mass–shift” property, which will play a crucial role not only in the present section but also in our later analysis of Hirzebruch surfaces, i.e., a ruled surface fibred over a rational curve.
@@
In the purely numerical optimisation below, we deliberately forget the dependence of $m=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}])$ on the full vector $t=(t_{1},\dots,t_{k})$ and retain only the coarse numerical constraints forced by the existence of the effective curves $E_{i}$ and $H-E_{i}-E_{j}$. Namely, for every Kähler class $\omega_{t}$ we have
@@
Suppose $\text{Bl}_{k}\mathbb{P}^{2}$ carries a PSC Kähler metric $\omega_{t}=\pi^{*}\omega-\sum t_{i}E_{i}$, where $\omega=aH$ is a Kähler metric on $\mathbb{P}^{2}$ for $a>0$. Then
$\mathcal{J}_{\text{Bl}_{k}\mathbb{P}^{2}}([\omega_{t}])<\mathcal{J}_{\mathbb{P}^{2}}([\omega])=12\pi.$
@@
are candidates for the holomorphic $2$-systole. Set $m:=\operatorname{sys}^{\operatorname{hol}}([\omega_{t}])\leq\min\{t_{i},\ a-t_{i}-t_{j}:\ 1\leq i\neq j\leq k\}$. Then
$\mathcal{J}_{\text{Bl}_{k}\mathbb{P}^{2}}([\omega_{t}])=4\pi\,m\cdot\frac{3a-S}{\,a^{2}-Q\,}=:4\pi\,\phi_{k}(t).$
We now analyze the supremum of $\phi_{k}$. Observe that $\mathcal{J}$ (and hence $\phi_{k}$) is invariant under rescaling of the Kähler class. Thus, without loss of generality, we may assume $a=1$ in what follows. With this normalization,
@@
By Proposition 3.1, we have
$Q\leq(k-1)m^{2}+\bigl(S-(k-1)m\bigr)^{2},$
with equality at
$t^{*}=\bigl(m,\dots,m,\,S-(k-1)m\bigr)$
up to permutation of coordinates. Since the denominator 1Q1-Q is decreasing in $Q$, we obtain
@@
whenever $t_{i}>0$, 3S>03-S>0 and 1Q>01-Q>0.
We now identify the $(m,S)$ for which the upper bound $F(m,S)$ is sharp, i.e. for which $t^{*}$ is still feasible and $m$ is still the minimum appearing in the definition of $\phi_{k}$. First, Proposition 3.1 requires $t_{i}^{*}\geq m$, i.e.
$S-(k-1)m\geq m\quad\Longleftrightarrow\quad S\geq km.$
Next, we need $m$ to remain the minimum in
$\{t_{i}^{*},\,1-t_{i}^{*}-t_{j}^{*}:\ 1\leq i\neq j\leq k\},$
so that $\phi_{k}(t^{*})$ is still of the form $m(3-S)/(1-Q(t^{*}))$. Since $t_{i}^{*}\geq m$, it suffices to ensure
$1-t_{i}^{*}-t_{j}^{*}\geq m\quad(1\leq i\neq j\leq k),$
which gives
$S\leq 1+(k-3)m.$
Finally, we must have $1-Q(t^{*})>0$, i.e.
$1-(k-1)m^{2}-\bigl(S-(k-1)m\bigr)^{2}>0.$
Collecting the constraints, the relevant domain for $(m,S)$ is
@@
For $(m,S)$ in this domain we have $\phi_{k}(t)\leq F(m,S)$, and equality holds for $t=t^{*}$. Now, introduce
$y:=S-(k-1)m,\qquad A:=1-(k-1)m^{2},\qquad c:=m\bigl(3-(k-1)m\bigr),$
so that $m\leq y\leq 1-2m$ and $A>0$ on the domain, and
$F(m,S)=\frac{c-my}{A-y^{2}}.$
Viewing $F$ as a function of $y$ (with $m$ fixed), a direct computation gives
$\frac{\partial}{\partial y}F(m,(k-1)m+y)=\frac{-mA-my^{2}+2cy}{(A-y^{2})^{2}}=:\frac{N(y)}{(A-y^{2})^{2}}.$
The numerator $N(y)$ is a concave quadratic in $y$ with a unique maximum at
$y_{0}=\frac{c}{m}=3-(k-1)m.$
One checks that $y_{0}\geq 1-2m$ for all admissible $m$, hence $N(y)\geq 0$ on [m,12m][m,1-2m]. Therefore, for fixed $m$, $F(m,S)$ is strictly increasing in $S$ on the domain (3.3), and
$\max_{\,km\leq S\leq 1+(k-3)m}F(m,S)=F\bigl(m,\,1+(k-3)m\bigr).$
@@
$F\bigl(m,\,1+(k-3)m\bigr)=\frac{m\bigl(2-(k-3)m\bigr)}{\,1-(k-1)m^{2}-(1-2m)^{2}\,}=\frac{2-(k-3)m}{\,4-(k+3)m\,}=:G_{k}(m),$ (3.4)
where the feasibility condition
$1-(k-1)m^{2}-(1-2m)^{2}>0$
is equivalent to
@@
$G_{k}^{\prime}(m)=\frac{-2k+18}{\bigl(4-(k+3)m\bigr)^{2}}.$
We now distinguish three cases.
Case 1: $2\leq k\leq 8$. Here 2k+18>0-2k+18>0, so $G_{k}$ is strictly increasing on $(0,4/(k+3))$. Since $4/(k+3)\geq 1/3$ for $k\leq 8$, the constraint $m\leq 1/3$ is active, and
@@
This value is attained at $m=\frac{1}{3}$, $S=\frac{k}{3}$, that is, $t_{1}=\dots=t_{k}=\frac{1}{3}$, and hence
$\sup_{\,t}\phi_{k}(t)=1\quad\text{for }2\leq k\leq 8.$
Case 2: $k=9$. In this case $G_{9}^{\prime}(m)\equiv 0$, so $G_{9}$ is constant on $(0,4/12)$. From (3.4) we obtain
$G_{9}(m)=\frac{2-6m}{4-12m}=\frac{1}{2},\qquad 0<m<\frac{1}{3}.$
Since $m<1/3$ (the endpoint $m=1/3$ is excluded by the strict inequality in the denominator condition), the value $G_{9}(m)=1/2$ is not attained but can be approached as $m\uparrow 1/3$, i.e. along vectors $t$ with $t_{i}\uparrow\frac{1}{3}$. Thus
$\sup_{\,t}\phi_{9}(t)=\frac{1}{2}.$
Case 3: $k\geq 10$. Now 2k+18<0-2k+18<0, so $G_{k}$ is strictly decreasing on $(0,4/(k+3))$ and
$\sup_{m}G_{k}(m)=\lim_{m\downarrow 0}G_{k}(m)=\frac{1}{2}.$
The corresponding extremal configurations are of the form
$t=(m,\dots,m,1-2m),\qquad m\downarrow 0,$
for which $m\to 0$ and $S\to 1$. Hence
$\sup_{\,t}\phi_{k}(t)=\frac{1}{2}\quad\text{for }k\geq 10.$
Collecting the three cases, we arrive at
$\sup_{\,t}\phi_{k}(t)=\begin{cases}1,&2\leq k\leq 8,\\[2.0pt] \frac{1}{2},&k\geq 9.\end{cases}$
@@
$\sup_{\,t}\phi_{k}(t)\leq 1\quad\text{for all }k\geq 2.$
Combining this with the case $k=1$ treated at the beginning, consequently, we obtain
@@
The dependence of the upper bound of $\mathcal{J}_{\text{Bl}_{k}\mathbb{P}^{2}}$ on $k$ is in perfect agreement with the standard classification of the blow-ups $\text{Bl}_{k}\mathbb{P}^{2}$ according to the positivity of the canonical class. Indeed,
$c_{1}^{2}(\text{Bl}_{k}\mathbb{P}^{2})=\left(3H-\sum_{i=1}^{k}E_{i}\right)^{2}=9-k,$
so that $\text{Bl}_{k}\mathbb{P}^{2}$ is del Pezzo for $k\leq 8$, satisfies $K_{\text{Bl}_{k}\mathbb{P}^{2}}$ nef and $c_{1}^{2}=0$ for $k=9$, and has $c_{1}^{2}<0$ for $k\geq 10$.
On the other hand, the algebraic optimization in the proof above leads to the one-variable function $G_{k}(m)$. The sign of $G_{k}^{\prime}(m)$ is exactly the sign of $c_{1}^{2}(X_{k})$. Consequently,
  • for $2\leq k\leq 8$ (del Pezzo, $c_{1}^{2}>0$) the worst-case upper bound for $\mathcal{J}_{X_{k}}$ is achieved at the largest admissible value of $m$;
  • for $k=9$ ($K_{X_{k}}$ nef, $c_{1}^{2}=0$) the function $G_{k}$ is flat, and the corresponding upper bound is approached by Kähler classes, but not attained;
  • for $k\geq 10$ ($c_{1}^{2}<0$) the worst case occurs in the opposite regime $m\to 0$.
We also note that for $k\geq 5$ additional (1)(-1)-curves appear, for instance conics of class $2H-\sum_{i\in I}E_{i}$ through five points. Their intersection with $\omega_{t}]$ gives further inequalities (such as $2a-\sum_{i\in I}t_{i}>0$) which can only shrink the feasible region for $(m,S)$ and hence potentially lower the true supremum of $\mathcal{J}_{X_{k}}$. Thus the upper bounds of $\mathcal{J}_{\text{Bl}_{k}\mathbb{P}^{2}}$ above should not be expected to be sharp in general. The fact that their kk–dependence mirrors the del Pezzo/nef/negative classification is a genuinely geometric phenomenon rather than an artefact of the estimates.
@@
with equality if and only if $k=0$, in which case $\omega_{0}$ is the Fubini-Study metric and $\operatorname{sys}_{2}(\omega)$ is achieved by $\mathbb{P}^{1}$. In particular, (3.5) holds strictly for all blow-ups $X_{k}$ with $k\geq 1$.
@@
For $k=2$, For $k=2$, any Kähler class can be written as $\omega]=aH-t_{1}E_{1}-t_{2}E_{2},\penalty 10000\ a>0,\ t_{i}>0$. And,
@@
and the first inequality is strict, since there is no cscK metric on $X_{k}$ for $k=1,2$.

### v2

@@ 3. The systolic inequality on $\mathbb{P}^{2}$ and its blow-up
$\mathcal{J}_{\mathbb{P}^{2}}([\omega])=4\pi\,\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\,\frac{c_{1}(\mathbb{P}^{2})\cdot[\omega]}{[\omega]^{2}}=4\pi\cdot a\cdot\frac{3a}{a^{2}}=12\pi.$ (3.1)
@@
for every Kähler class $\omega]$ on $X_{k}$ and every $k\geq 1$. Since $\mathcal{J}_{X_{k}}$ is defined purely in terms of intersection numbers, the problem is intrinsically finite-dimensional. However, the complexity of the Mori cone $\overline{\mathrm{NE}}(X_{k})$ increases rapidly with $k$: for large $k$ there are many extremal rays, and for $k\geq 9$ (in the non–del Pezzo regime) the cone is not even finitely generated. A direct ray–by–ray analysis of all effective curve classes is therefore hopeless.
@@
For a given Kähler class we set $m:=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$. In particular, since each exceptional curve $E_{i}$ is effective and $\omega]\cdot E_{i}=t_{i}$, we always have $m\leq t_{i}$ for all $i$. Thus, if we fix $(a,m,S)$ and vary the individual $t_{i}$ subject to
@@
then $a$ and $S$ are held fixed while $\mathcal{J}_{X_{k}}([\omega])$ changes only through the quantity $Q$ appearing in $\omega]^{2}$. In particular, for fixed $(a,m,S)$ the maximal possible value of $\mathcal{J}_{X_{k}}([\omega])$ is attained when $Q$ is as large as allowed by these constraints. This leads to the following elementary “mass–shift” property, which will play a crucial role not only in the present section but also in our later analysis of Hirzebruch surfaces, i.e. geometrically ruled surfaces over $\mathbb{P}^{1}$.
@@
In the numerical optimization below, we enlarge the geometric problem to a coarser numerical one. More precisely, we deliberately forget the full dependence of $m=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}])$ on the vector $t=(t_{1},\dots,t_{k})$, and retain only the necessary inequalities forced by the effective curves $E_{i}$ and $H-E_{i}-E_{j}$. This produces a larger numerical feasible region, so any upper bound obtained there automatically yields an upper bound for the geometric supremum of $\mathcal{J}_{X_{k}}$. Namely, for every Kähler class $\omega_{t}$ we have
@@
Let $X_{k}:=\operatorname{Bl}_{k}\mathbb{P}^{2}$, and let $\omega_{t}]=aH-\sum_{i=1}^{k}t_{i}E_{i}\in\mathcal{K}^{+}(X_{k})$, where $a>0$ and $k\geq 1$. Then
$\mathcal{J}_{X_{k}}([\omega_{t}])<\mathcal{J}_{\mathbb{P}^{2}}(aH)=12\pi.$
@@
are candidates for the holomorphic $2$-systole. Set $m:=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}])\leq\min\{t_{i},\ a-t_{i}-t_{j}:\ 1\leq i\neq j\leq k\}$. Then
$\mathcal{J}_{X_{k}}([\omega_{t}])=4\pi\,m\,\frac{3a-S}{a^{2}-Q}.$
For convenience, we denote the corresponding numerical factor by
$\phi_{k}(t):=m\,\frac{3a-S}{a^{2}-Q},$
where $m=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}])$, $S=\sum_{i}t_{i}$, and $Q=\sum_{i}t_{i}^{2}$. We now estimate $\phi_{k}(t)$ from above by optimizing over the enlarged numerical feasible region described above. Observe that $\mathcal{J}$ (and hence $\phi_{k}$) is invariant under rescaling of the Kähler class. Thus, without loss of generality, we may assume $a=1$ in what follows. With this normalization,
@@
By Proposition 3.1, we have $Q\leq(k-1)m^{2}+\bigl(S-(k-1)m\bigr)^{2}$, with equality at $t^{*}=(m,\dots,m,\,S-(k-1)m)$ up to permutation of coordinates. Since the denominator 1Q1-Q is decreasing in $Q$, we obtain
@@
for every geometric vector $t$ arising from a class in $\mathcal{K}^{+}(X_{k})$. We now work on the admissible domain
@@
We distinguish two parts.
Case 1: $2\leq k\leq 8$. Set B:=1+(k3)mB:=1+(k-3)m. We claim that for every $(m,S)$ in the admissible domain, $F(m,S)\leq F(m,B)$. Indeed, a direct computation gives
$F(m,B)-F(m,S)=\frac{(B-S)\,L_{k}(m,S)}{\bigl(4-(k+3)m\bigr)\bigl(1-(k-1)m^{2}-(S-(k-1)m)^{2}\bigr)},$
where $L_{k}(m,S)=S(2-(k-3)m)+k(k-1)m^{2}-3(k+1)m+2$.
Since $B-S\geq 0$, 4(k+3)m>04-(k+3)m>0, and $1-(k-1)m^{2}-(S-(k-1)m)^{2}>0$, it remains to show that $L_{k}(m,S)\geq 0$. For $2\leq k\leq 8$ and $0<m\leq 1/3$, we have
$2-(k-3)m\geq 2-\frac{k-3}{3}=\frac{9-k}{3}>0,$
so $L_{k}(m,S)$ is increasing in $S$. Hence
$L_{k}(m,S)\geq L_{k}(m,km)=2k\,m^{2}-(k+3)m+2.$
The discriminant of the quadratic polynomial $2k\,m^{2}-(k+3)m+2$ is
$(k+3)^{2}-16k=(k-1)(k-9)<0\qquad(2\leq k\leq 8),$
and its leading coefficient is positive. Therefore $2k\,m^{2}-(k+3)m+2>0$ for all $m$, and hence $L_{k}(m,S)>0$. This proves
$F(m,S)\leq F\bigl(m,1+(k-3)m\bigr).$
@@
$F\bigl(m,1+(k-3)m\bigr)=\frac{2-(k-3)m}{\,4-(k+3)m\,}=:G_{k}(m).$ (3.4)
The feasibility condition $1-(k-1)m^{2}-(1-2m)^{2}>0$ is equivalent to
@@
$G_{k}^{\prime}(m)=\frac{18-2k}{(4-(k+3)m)^{2}}>0\qquad(2\leq k\leq 8).$
Since $4/(k+3)\geq 1/3$ for $k\leq 8$, the constraint $m\leq 1/3$ is active, and thus
@@
This value is attained at
$m=\frac{1}{3},\qquad S=1+(k-3)\frac{1}{3}=\frac{k}{3}.$
Hence
$\sup_{t}\phi_{k}(t)\leq 1\qquad\text{for }2\leq k\leq 8.$
Case 2: $k\geq 9$. We claim that
$F(m,S)\leq\frac{1}{2}$
throughout the admissible domain. Indeed, this inequality is equivalent to
$2m(3-S)\leq 1-(k-1)m^{2}-(S-(k-1)m)^{2}.$
Set y:=S(k1)my:=S-(k-1)m. Then the difference between the right-hand side and the left-hand side is
$1-(k-1)m^{2}-y^{2}-2m(3-(k-1)m-y)=1-(y-m)^{2}+km^{2}-6m.$
Since $km\leq S\leq 1+(k-3)m$, we have $m\leq y\leq 1-2m$, and therefore $0\leq y-m\leq 1-3m$. It follows that
$1-(y-m)^{2}+km^{2}-6m\geq 1-(1-3m)^{2}+km^{2}-6m=(k-9)m^{2}\geq 0.$
Thus
$F(m,S)\leq\frac{1}{2}\qquad\text{for all }k\geq 9.$
For $k=9$, equality is achieved whenever $S=1+6m$, equivalently y=12my=1-2m, so
$\sup_{t}\phi_{9}(t)\leq\frac{1}{2}.$
[…]
$\sup_{t}\phi_{k}(t)\leq\frac{1}{2}\qquad\text{for }k\geq 10.$
Collecting the two cases, we conclude that
$\sup_{t}\phi_{k}(t)\leq\begin{cases}1,&2\leq k\leq 8,\\[2.0pt] \frac{1}{2},&k\geq 9.\end{cases}$
@@
$\sup_{t}\phi_{k}(t)\leq 1\qquad\text{for all }k\geq 2.$
Combining this with the case $k=1$ treated at the beginning, we obtain
@@
The dependence of the upper bounds for $\mathcal{J}_{\operatorname{Bl}_{k}\mathbb{P}^{2}}$ on $k$ is in perfect agreement with the standard classification of the blow-ups $\operatorname{Bl}_{k}\mathbb{P}^{2}$ according to the positivity of the canonical class. Indeed,
$c_{1}^{2}(\operatorname{Bl}_{k}\mathbb{P}^{2})=\left(3H-\sum_{i=1}^{k}E_{i}\right)^{2}=9-k.$
Thus $k=9$ is exactly the borderline case where $c_{1}^{2}=0$, while $c_{1}^{2}>0$ for $k\leq 8$ and $c_{1}^{2}<0$ for $k\geq 10$. In the del Pezzo range $k\leq 8$ (for points in general position), this agrees with the usual positivity picture. We also note that for $k\geq 5$ additional (1)(-1)-curves appear, for instance conics of class $2H-\sum_{i\in I}E_{i}$ through five points. Their intersection with $\omega_{t}]$ gives further inequalities (such as $2a-\sum_{i\in I}t_{i}>0$) which can only shrink the feasible region for $(m,S)$ and hence potentially lower the true supremum of $\mathcal{J}_{X_{k}}$. Thus the upper bounds of $\mathcal{J}_{\text{Bl}_{k}\mathbb{P}^{2}}$ above should not be expected to be sharp in general.
@@
with equality if and only if $k=0$, in which case $\omega_{0}$ is the Fubini-Study metric (after normalization) and $\operatorname{sys}_{2}(\omega)$ is achieved by $\mathbb{P}^{1}$. In particular, (3.5) holds strictly for all blow-ups $X_{k}$ with $k\geq 1$.
Proof.
For $k=0$, equation (3.1) shows that $\mathcal{J}_{\mathbb{P}^{2}}([\omega])=12\pi$ for every Kähler class $\omega]$ on $\mathbb{P}^{2}$. Hence Theorem 2.2 (1) gives
$\min_{\mathbb{P}^{2}}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 12\pi.$
For $k\geq 1$, Lemma 3.2 shows that every class $\omega_{k}]\in\mathcal{K}^{+}(X_{k})$ satisfies
$\mathcal{J}_{X_{k}}([\omega_{k}])<12\pi.$
If $\omega_{k}$ is a PSC Kähler metric on $X_{k}$, then $\omega_{k}]\in\mathcal{K}^{+}(X_{k})$, so Theorem 2.2 (1) again yields
$\min_{X_{k}}S(\omega_{k})\cdot\operatorname{sys}_{2}(\omega_{k})\leq\mathcal{J}_{X_{k}}([\omega_{k}])<12\pi.$
This proves the strict inequality for all $k\geq 1$.
It remains to characterize the equality case for $k=0$. If
$\min_{\mathbb{P}^{2}}S(\omega_{0})\cdot\operatorname{sys}_{2}(\omega_{0})=12\pi,$
then equality must hold in Theorem 2.2 (1). Hence $\omega_{0}$ is cscK and $\operatorname{sys}_{2}(\omega_{0})$ is realized by a holomorphic $1$-cycle. Since every Kähler class on $\mathbb{P}^{2}$ is proportional to $H$, such a cscK metric is homothetic to the Fubini–Study metric, and the holomorphic $2$-systole is realized by a projective line $\mathbb{P}^{1}\subset\mathbb{P}^{2}$. Conversely, every metric homothetic to the Fubini–Study metric realizes equality. ∎
@@
For $k=2$, any Kähler class can be written as $\omega]=aH-t_{1}E_{1}-t_{2}E_{2},\penalty 10000\ a>0,\ t_{i}>0$. And,
@@
and the inequality is strict since there is no cscK metric on $\operatorname{Bl}_{k}\mathbb{P}^{2}$ for $k=1,2$.